A Complete Bibliography of Publications in Mathematics and Computers in Simulation: 1960–1969

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.00

Title word cross-reference

e [364].

000 [475].

10.000 [475]. 110 [461]. 1960 [11, 1, 22, 35]. 1961 [58, 47, 69, 80]. 1962 [104, 92, 118, 130]. 1963 [153, 143, 170, 184]. 1964 [208, 195, 221, 233, 300]. 1965 [258, 245, 269, 279]. 1966 [301, 291, 312, 322]. 1967 [353, 335, 363, 374]. 1968 [403, 388, 415, 427, 439]. 1969 [448, 438, 458, 470].

a-c-network [40]. A.D.C. [251]. Activité [20, 33, 9, 57, 68, 79, 90, 103, 116, 128, 140, 194, 151, 168, 182, 220, 207, 232, 243, 287, 257, 268, 278, 333, 299, 311, 321, 352, 362, 373, 387, 426, 402, 414, 436, 447, 457, 469, 482, 44]. ADAC [339]. additive [452]. Adjoint [134]. Adjusting [316]. aerodynamic [451]. aeronautique [185]. AICA [391]. aid [62, 132]. aide [365, 284]. algebraic [262, 51]. allocation [98]. alloy [340]. amplifier [108, 324, 74]. amplifiers [409, 120]. amplitude [418]. analac [224, 461]. Analog [357, 161, 463, 306, 197, 356, 429, 209, 62, 98, 461, 260, 249, 380, 355, 210, 393, 381, 188, 164, 50, 73, 85, 84, 211, 261, 419, 315, 280, 93, 134, 163, 97, 38, 406, 417, 283, 471, 451, 198, 314, 378, 23, 24, 175, 225, 337, 305, 339, 132, 201, 252, 476, 171, 147, 338, 439, 236, 95]. analog-computer [210, 93, 163]. analog-digital [378, 226, 200, 476, 147]. analog-to-digital [161]. Analogie [14]. analogique [461, 450, 365, 234, 223, 59, 274, 105, 284, 185, 27, 71, 315, 463, 24, 337, 270, 375, 472, 431, 405, 16, 273, 37, 49, 94, 293, 171, 147, 439, 95]. analogique-digitale [147]. analogiques [280, 451]. Analogrechner [406, 164, 85, 417, 338, 97]. Analogrechners [73, 38]. Analogue [300, 60, 70, 106, 235, 36, 212, 290, 110, 72, 251, 340, 157, 122, 172, 48]. analogy [14]. analyser [96, 187, 28, 51, 202]. analysis [347, 5, 345, 346, 344, 198, 342, 197, 356, 201]. analytic [262]. Analytical [28]. analyzer [294, 145, 174, 40]. analyzers [63]. angle [418]. annals [290]. antiseismic [280]. antiseismique [280]. Apache [157]. APE [393]. appearance [390]. Application [201, 147, 174]. applications [237, 16]. applied [60, 106, 439]. appliqué [439]. appliquees [105]. approach [380, 409, 262]. April [11, 58, 104, 153, 208, 258, 301, 353, 403, 448]. arbitrary [112, 368]. arithmetic [262]. Arrangement [227]. aseismic [367, 382]. asservissements [273]. associated [14, 451]. association [20, 33, 194, 220, 287, 290, 333, 352, 426]. associée [14]. associées [451]. ASTRAC [210, 271, 199, 378]. ASTRAC-I [271]. asymmetric [366]. atmospheric [109]. atomic [60]. atomique [59]. Authors [395, 290]. auto [178]. Automated [343, 440]. Automatic [249, 176, 380, 70, 250, 158, 119]. automatically [460]. automatique [473, 274, 71, 176]. Automatische [119]. automobile [306]. autour [337, 39]. aux [147]. avenir [94]. averaging [378]. axial [83, 14]. axiale [14].

Backlash [72]. band [340, 108, 160]. beat [407]. Behandlung [97]. behavior [355, 73]. behaviour [408]. Berkeley [162]. besoins [59]. Betriebsverhaltens [73]. between [172, 248, 171]. bibliographique [7, 18, 31, 42, 54, 65, 76, 87, 100, 114, 126, 138, 149, 166, 180, 191, 205, 218, 230, 241, 255, 266, 276, 286, 296, 308, 318, 330, 349, 359, 370, 384, 399, 411, 423, 433, 444, 454, 466, 479]. borders [451]. boundaries [366]. boundary [441, 263]. boundary-value [441]. Brighton [300]. budget [50]. Busch [15, 15].

c [5, 40]. C.H.R. [475]. cable [357]. Calcul [71, 379, 234, 59, 274, 105, 185, 24, 270, 375, 16, 323, 37, 49, 94, 171, 439]. calculateur [461, 365, 224, 223, 284, 315, 337, 475]. calculateurs [280]. calculating [379]. calculation [24, 171, 439]. calculator [224, 164, 73, 85, 315, 38, 406, 337, 475, 338]. calculators [97]. calculatrice [431, 273]. calibration [160]. carrier [74]. cases [164]. ce [270]. centre [39]. certain [164]. champ [284, 337]. champs [405]. Changement [274]. character [248]. charged [176]. chargées [176]. checking [419]. choix [354]. Chronique [7, 18, 31, 42, 54, 65, 76, 87, 100, 114, 126, 138, 149, 166, 180, 191, 205, 218, 230, 241, 255, 266, 276, 286, 296, 308, 318, 330, 349, 359, 370, 384, 399, 411, 423, 433, 444, 454, 466, 479]. Circuit [123, 251, 340, 135]. circuits [110]. civil [36, 37]. classical [452]. closer [172]. code [157, 294]. coherent [452]. column [248]. Combined [225, 476]. commandabilite [472]. commande [473]. Comment [226]. communication [119]. communications [271]. commutation [475, 293]. company [15]. comparator [161]. Comparison [61]. complex [97]. comportant [451]. Computation [300, 290, 60, 106, 419, 134, 343, 23, 175, 225, 36, 476, 48, 159, 172]. computations [211]. computer [62, 98, 461, 260, 249, 355, 210, 392, 393, 357, 381, 188, 340, 345, 50, 157, 261, 15, 93, 163, 417, 283, 471, 122, 378, 306, 158, 339, 132, 429, 201, 252, 212, 236]. computers [380, 110, 72, 84, 119, 246, 280, 225, 209]. computing [70, 420, 343, 74]. computing-error [343]. computor [304, 418, 305]. concept [212]. conclusions [391]. condensation [405]. condition [147]. conduction [164, 145]. conformal [51]. Congrès [364]. considérations [379]. constraints [250]. construction [315, 431, 405]. contact [172, 171]. containing [198, 314]. Contribution [273]. control [249, 70, 393, 441, 250, 215, 158, 197, 449]. controles [273]. controllers [407]. conversion [161]. correction [419]. correlation [122, 178]. Correspondance [377, 394, 376]. cost [84]. coupled [107]. courants [461]. courbe [365]. critere [354]. criteria [198]. criterion [197]. cross [178]. cross-correlation [178]. cumulative [187]. current [40]. curve [328]. curves [238]. cylindres [473]. Cylindrical [381].

d [5]. d-c [5]. d.c [357]. D.C. [72]. d’aèrodynamique [451]. dans [365, 27, 431, 405, 39, 16]. d’articles [443]. d’articlse [8, 19, 32, 43, 55, 66, 77, 88, 101, 115, 127, 139, 150, 167, 181, 192, 203, 216, 228, 239, 253, 264, 272, 285, 295, 307, 317, 329, 348, 358, 369, 383, 398, 410, 422, 432, 453, 465, 477]. data [471, 337]. dead [407]. d’echanges [462]. d’echelle [274]. decisions [161]. d’ecoulements [284]. defined [213]. delay [135, 212]. delays [476]. densite [27]. depend [354]. dependant [212]. dependent [417]. d’équations [463]. descent [326]. Design [407, 367, 382, 198]. detection [452, 452]. Determination [263, 472, 188, 83, 73]. Determining [248]. devant [59]. Development [214]. developments [23]. developpements [146]. deviation [357]. d’exploitation [396]. Diffe [460]. diffential [62]. difference [107]. differences [198]. Differential [96, 392, 186, 344, 97, 283, 294, 174, 28, 51, 63, 460, 236]. differential-equation-solving [460]. differentialgleichungen [97]. difficulties [164]. diffusion [281, 107, 164, 463, 405]. Diffusions [164]. Diffusions-und [164]. Digital [199, 393, 157, 407, 161, 451, 378, 226, 200, 63, 476, 147]. digitale [147]. digitales [451]. dimensional [366]. d’impedances [146]. d’impulsions [273]. d’integration [29]. d’inversion [16]. diode [461, 52, 252, 247]. diodes [397]. Dirchlet [237]. direct [339]. Dirichlet [237, 337]. discontinuous [135]. discrete [123]. display [460]. distortion [120]. distributed [341]. distribution [234, 187]. division [121]. Documents [10, 21, 34, 45, 56, 67, 78, 89, 102, 117, 129, 141, 152, 169, 183, 193, 206, 219, 231, 242, 256, 267, 277, 288, 297, 309, 319, 331, 350, 360, 371, 385, 400, 412, 424, 434, 445, 455, 467, 480]. domain [14]. domaine [14]. donneés [337]. d’optimalisation [450]. double [162]. due [347, 357, 108]. d’un [284, 14, 315, 451, 337, 147]. d’une [461, 365, 27, 431]. Dynamic [367, 382, 159, 355, 73, 476, 236]. dynamischen [73].

Easecomputer [162]. échanges [95]. echantillonnes [472]. écoulement [337]. editor [173, 156]. Editorial [13, 404, 12, 131, 428, 154, 82, 222, 302, 303, 155, 196, 313, 292, 81, 259, 336]. education [209, 48]. effects [252, 147]. effets [147]. eigenvalues [188]. einem [406, 417]. electrical [73, 197]. electrique [234]. electromagnetic [83]. electromechanical [85]. electronic [316, 72, 251, 73, 85, 280, 97, 38, 406, 417, 475, 338]. électronique [475]. électroniques [280]. elektrischer [73]. elektromechanischen [85].

Elektronischen [417, 73, 85, 97, 38, 406]. elektronischer [338]. élémentaires [147]. elementary [147]. elements [405]. elevated [367, 382]. énergétiques [379]. energie [234, 439]. energy [439]. engineering [260, 60, 36, 209, 430]. Enseignement [49, 105]. entrainees [365]. entre [284, 171]. equation [365, 281, 164, 186, 460, 236]. equations [62, 392, 381, 85, 344, 97, 283, 463, 28, 227, 124]. equivalent [340]. Ermittlung [73]. Errata [91, 478]. Error [5, 345, 344, 346, 163, 343, 342, 198]. Error-analysis [345]. errors [347, 108, 476, 236]. étroit [171]. Etude [284, 27, 365, 405, 273, 461, 315, 337]. evaluating [162]. evaluation [236, 237]. Exact [186]. exchanger [392, 61]. exchanges [95]. Exemple [450]. expandable [50]. expansion [199]. Experimental [160, 16]. exponentielle [325]. extension [135].

Fälle [164]. fast [343]. feedback [249, 197]. ferromagnetic [355]. ferroresonant [251, 356]. field [144, 337]. fields [366]. fifth [391]. films [355]. filters [160]. Finite [198, 107]. Firma [15]. first [186, 283, 236]. first-order [283, 236]. flow [337, 145]. flowing [198, 314]. fluids [198, 314]. flutter [174]. flux [365]. follower [328]. fonction [325, 14]. Fonctions [133]. Force [379]. forcée [379]. forces [83]. formulation [442]. four [405]. Fourier [124]. Fourier-type [124]. Fourth [300]. fractionnee [29]. fre [136]. free [451]. French [379, 461, 224, 2, 237, 464, 14, 315, 280, 463, 451, 24, 337, 133, 475, 176, 171, 147, 439, 95]. frequency [74]. frontières [451]. function [271, 4, 14, 327, 212, 247]. functions [112, 304, 418, 326, 235, 133, 52, 368, 178]. future [93].

gain [324]. Gaussian [452]. General [379, 380]. générale [379]. generalised [212]. Generateur [325]. Generation [112, 418, 326, 136, 368]. generator [4, 213, 163, 327, 52, 247]. genie [59, 37]. German [164, 73, 85, 119, 15, 97, 38, 406, 417, 338]. Gewisser [164]. gradient [406, 429]. Gradientenmethode [406]. grande [431]. gravite [39].

harmonic [14]. harmonique [14]. heat [392, 164, 61, 440, 145, 248]. heterogeneous [95]. hétérogènes [95]. high [210]. high-speed [210]. Higher [48]. Hilfe [73, 406]. Hofmann [15]. Hybrid [392, 474, 420, 378, 380, 282, 345, 419, 346, 344, 343, 226, 200, 158, 294, 475, 476, 159, 439, 214]. hybrid-code [294]. hybrid/analog [226, 200]. hybrid/analog-digital [226, 200]. hybride [323, 475, 439]. hydrometallurgique [27]. hysteresis [356].

Identification [459, 342]. II [393, 378]. illimité [14]. im [97]. immittance [409]. impulse [160]. including [356]. Index [46, 142, 244, 334, 437]. induction [474]. industrie [234]. Informations [6, 17, 30, 41, 53, 64, 75, 86, 99, 113, 125, 148, 165, 179, 190, 204, 217, 229, 240, 254, 265, 275, 289, 298, 310, 320, 332, 351, 361, 372, 386, 401, 413, 425, 435, 446, 456, 468, 481, 137]. inherent [476]. input [248]. inputs [202]. integral [237]. intégrale [237]. integrals [162]. integrating [282]. integration [84, 39]. integrator [5]. intended [157]. International [300, 290, 439, 439]. Interrelation [342]. interrupteur [315]. introduced [476]. inventer [474]. inventer-induction [474]. inverse [326]. invertor [409]. investigation [366]. Iterative [211, 314, 223, 393, 261].

January

[1, 47, 92, 143, 195, 245, 291, 335, 388, 438]. jet [215]. July [22, 69, 118, 170, 221, 269, 312, 363, 415, 458]. junction [340].

kind [328]. Kommunikation [119]. komplexen [97].

laboratoire [94]. laboratory [93]. l’AICA [364]. laminoirs [473]. Laplacian [337]. laplacien [337]. largeur [273]. l’Association [9, 57, 68, 79, 90, 103, 116, 128, 140, 151, 168, 182, 207, 232, 243, 257, 268, 278, 299, 311, 321, 362, 373, 387, 402, 414, 436, 447, 457, 469, 482, 44]. Letter [173, 156, 395]. level [441]. level-control [441]. LHB [15, 15]. libres [451]. Liebmann [440]. limitations [108]. limites [147]. limits [147]. line [471, 460]. linéaire [451, 147]. linéaires [379, 463, 133]. linear [379, 249, 134, 306, 135]. lines [461]. Linke [15, 15]. Linke-Hofmann-Busch [15, 15]. logic [430]. look [134]. loop [356]. loops [345]. Losung [38, 164, 406, 417]. Low [84, 136]. Low-cost [84]. low-fre [136].

machines [73, 262]. made [247]. magnetique [284]. magnetiques [293]. mapping [51]. Maschinen [73]. mass [145]. mathematical [366]. mathematics [106, 246, 132]. mathematiques [105]. Matric [246, 304, 418, 262]. matrices [188]. means [83, 157, 40]. measurements [177]. meeting [391]. Meetings [300]. memory [110, 175, 421]. memory-pair [421]. Mesure [237]. method [379, 188, 441, 419, 407, 406, 122, 326, 51, 160]. Methode [29, 379]. methods [314]. minimum [408]. Mitrovic [197]. Mittels [38]. model [474, 366, 197]. models [198]. Modern [393]. modes [263]. modulaire [462]. modulation [418, 273]. modulator [420]. module [393]. money [98]. motor [474]. mouvement [365, 39]. Mr. [395]. multi [161]. multi-comparator [161]. multidimensionnels [442]. multiple [161]. multiplicateur [397]. multiplier [121, 327]. multipliers [84]. Multiplying [282]. mutually [107].

Nachbildung [85]. needs [60]. network [83, 15, 202, 40]. networks [440, 408]. Netzwerkrechenmaschine [15]. noise [452, 122]. non [463, 451, 306, 408, 133, 202, 147]. non-linear [306]. non-minimum [408]. non-sinusoidal [202]. Nonlinear [133, 463, 451, 147]. normal [263]. Nouveaux [365]. nouvelle [389]. Nouvelles [16]. novel [327]. noyaux [293]. nucléaire [439]. nucléaires [95]. nuclear [439, 95].

observabilite [472]. obtaining [122].

obtenus [365]. October

[35, 80, 130, 184, 233, 279, 322, 374, 427, 470]. On-line [471, 460]. one [226, 200]. one-parameter [226, 200]. operated [421]. operating [73]. operational [108, 409, 324, 120]. operator [107]. Optimal [368, 449]. Optimalisation [223]. Optimierungsaufgaben [406, 417]. optimisation [464]. optimization [249, 464, 163, 406, 417, 429]. optimizer [226, 200]. Optimum [189]. order [186, 283, 236, 449]. originales [431]. orthogonal [271]. orthogonal-function [271]. oscillation [379]. oscillator [111, 409].

pace [122]. Pacit [50]. Pages

[11, 1, 22, 35, 58, 47, 69, 80, 104, 92, 118, 130, 153, 143, 170, 184, 208, 195, 221, 233, 258, 245, 269, 279, 301, 291, 312, 322, 353, 335, 363, 374, 403, 388, 415, 427, 448, 438, 458, 470]. pair [421]. paper [327]. paper-tape [327]. parallel [430]. paralleles [284]. Parameter [261, 163, 341, 342, 226, 200, 429]. Parameter-perturbation [163].

Parametern [417]. parameters [417]. parasitic [252]. partial [392, 186, 344, 283, 236]. particles [176]. particules [365, 176]. partir [379]. pass [108, 160]. performance [210]. perturbation [163]. Phase [120, 111, 452, 408]. phase-coherent [452]. phase-shift [408]. physical [198, 314]. plan [337]. planar [337]. plans [284]. plants [449]. plus [171]. plusieurs [133]. point [441, 263]. polynomials [51]. porteurs [461]. positionnement [473]. potential [14]. potentiel [14]. Pour [171, 315]. power [461, 40]. precision [50]. Preliminary [328]. presence [284]. Présentation [389]. principe [16]. probability [187]. problem [452, 96, 263, 157, 451, 174, 109, 147]. probleme [450, 451, 147]. problems [98, 441, 38, 440, 40]. Procede [396]. process [38]. processes [107]. Processing [178]. processus [27]. production [234]. professionals [171]. professionnels [171]. profil [337]. profile [337]. profiles [189]. program [214]. programmed [339]. programming [157, 159]. Progrès [24]. progress [24]. properties [213]. propos [26]. provided [147]. publication [290]. pulse [452, 420]. purpose [380].

que [270]. Quelques [431, 39, 461]. quency [136]. Qu’est [270]. Qu’est-ce [270].

random [213]. rapide [396]. rate [347]. ratio [420, 248]. RC [409]. réacteurs [95]. reaction [215]. reactor [214]. reactors [63, 189, 95]. real [464]. réalisation [315]. realization [409]. receiver [177]. recents [146, 24]. Rechenanlagen [119]. recherche [280]. Recommendations [290]. recorder [177]. rectification [248]. recus [10, 21, 45, 56, 67, 78, 89, 102, 117, 129, 141, 152, 169, 183, 193, 206, 231, 242, 256, 267, 277, 288, 297, 309, 319, 331, 350, 360, 371, 385, 400, 412, 434, 445, 455, 467, 480, 34, 219, 424]. redresseurs [273]. reduction [471]. redundant [419]. réel [464]. Reflexions [323]. reflux [248]. regime [463, 463]. regimes [461, 461]. réglage [71]. regulation [27]. relation [248]. Relay [421]. remarques [39]. remote [471]. repetitive [225, 174, 51]. Replica [85]. reply [395]. report [328]. Representation [405, 293, 95]. rescaling [380]. research [280]. reseaux [281, 396, 475, 146]. resistance [83]. resistances [396]. resistifs [281]. resistors [316]. Resolution [281, 450]. Resonance [188]. response [251, 178]. resultats [365]. Results [391]. Résumés [8, 19, 32, 43, 55, 66, 77, 88, 101, 115, 127, 139, 150, 167, 181, 192, 203, 216, 228, 239, 253, 264, 272, 285, 295, 307, 317, 329, 348, 358, 369, 383, 398, 410, 422, 432, 443, 453, 465, 477]. review [390]. revue [389]. rf [452]. Rheoelectric [14]. rhéoélectrique [14]. roulette [238]. roulette-type [238]. routine [460].

Salzgitter [15]. Salzgitter-Watenstedt [15]. sample [213, 378]. sample-averaging [378]. sample-time [213]. sampling [347]. saturable [63]. scaled [460]. scaling [227]. Schalter [338]. Schwingungsproblemen [38]. science [430]. sein [95]. sensitivity [342]. September [300, 439]. septembre [439]. sequentially [324]. sequentially-switched [324]. series [251]. ses [237]. several [133]. sharing [175]. shift [408]. ship [357]. Shock [124]. Siderurgie [375]. simple [397, 61]. Simrar [177]. Simulated [177]. simulateur [462, 39]. Simulating [98, 61]. Simulation [452, 473, 442, 215, 463, 341, 109, 72, 251, 340, 186, 144, 314, 235, 459, 356, 63, 212, 214, 416]. simulator [145]. simulators [135]. sinusoidal [202]. situation [124]. slow [225, 305]. slow-type [225]. small [393, 50]. small-budget [50]. solid [121]. solide [365]. solidifying [96]. solution [392, 357, 164, 38, 406, 417, 283, 440, 28, 147, 236]. solutions [344, 262, 431]. Solving [441, 40, 460]. Some [238, 461]. sommaire [244, 437, 46, 142, 334]. sources [122, 40]. space [215]. special [186]. specialists [172]. Spectral [346]. speed [210]. speeds [252]. Stabilitätsschwierigkeiten [164]. stabilite [284]. stability [262]. stabilizer [164]. state [121, 108]. static [236]. stationary [147]. stationnaire [147]. statistical [177, 213]. steady [108]. steady-state [108]. steel [96]. steepest [326]. step [178]. stressed [355]. studies [163, 343]. Study [461, 305, 476, 355, 210, 306, 145, 252, 271, 315, 337]. subharmonic [251]. submarine [357]. subroutines [346]. suitable [247]. summation [123]. supérieur [49]. surface [305]. suspension [365, 306]. sweep [261]. switch [315, 338]. switched [324]. switching [475, 252]. symétrie [14]. symmetry [83, 14]. Symposium [439]. synthesis [250]. system [379, 210, 271, 474, 199, 85, 215, 306, 356, 40]. systemes [442, 464, 379]. Systemgleichungen [85]. systems [260, 249, 250, 464, 134, 262, 97, 198, 314, 341, 472, 197, 476].

tape [327]. tasks [406, 417]. teaching [62, 260, 132, 430]. technical [3, 25]. technique [26, 38, 202]. techniques [261, 451, 341, 429]. temperature [189]. temps [464]. Tendances [146]. tensile [471]. test [471]. texts [290]. theorem [135]. theory [246, 135]. thermal [381, 95]. thermiques [462, 405, 95]. thin [355]. third [449]. Three [111, 366]. three-dimensional [366]. Time [175, 449, 112, 464, 121, 213, 417, 212]. time-dependant [212]. time-dependent [417]. time-division [121]. Time-sharing [175]. tower [367, 382]. tracer [176]. Traceur [176]. train [452]. training [106].

Traitement [451]. trajectoires [176].

trajectories [176]. Transfer

[304, 440, 235, 212]. transformer [83]. Transient [408]. transistor [340]. transistorized [74]. transistors [315]. transitional [463]. transitoire [463]. transitoires [462]. transitores [461]. transitory [461]. transport [234, 109]. Treatment [451, 96, 97]. Trends [144]. triangle [84]. triangle-integration [84]. triangular [136]. Trigonometric [51]. tubular [61, 189, 214]. tunnel [461, 252]. tunnel-diode [252]. two [441, 263, 4, 61, 52, 247]. two-point [441, 263]. two-variables [4]. type [238, 225, 124].

units [420]. universal [224]. universel [224]. unknown [452]. unlimited [14]. Use [162, 260, 110, 197, 28, 51, 430, 280]. using [73, 409, 420, 38, 406, 471, 212]. Utilisation [280]. utilisees [431]. utilizing [175].

V [364]. value [441, 263]. variable [324, 122, 337]. variable-gain [324]. variables [4, 419, 133, 52, 247]. variation [347]. varying [449]. vehicle [215]. Verfahrenstecnik [38]. Versailles [439]. vibration [38]. VII [334]. vocabulaire [2, 26]. vocabulary [3, 2, 25]. voie [450]. voltages [136]. volume [334]. Vroman [395].

Wärmeleitungsgleichung [164].

Watenstedt [15]. water [367, 382]. wave [305]. waveform [201]. way [282, 316]. ways [61]. wide [340]. windings [83]. within [95].

Zeitabhaengigen [417]. Zur [38].

References

Anonymous:1960:PJa

[1] Anonymous. Pages 2–51 (January 1960). Mathematics and Computers in Simulation, 2(1):??, January 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Danloux-Dumesnils:1960:DVF

[2] M. Danloux-Dumesnils. Du vocabulaire. (French) [On vocabulary]. Mathematics and Computers in Simulation, 2(1):2–3, January 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800018.

Anonymous:1960:ATV

[3] Anonymous. About technical vocabulary. Mathematics and Computers in Simulation, 2(1):4, January 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546080002X.

Gonzalez-Lbeas:1960:TVF

[4] J. Gonzalez-Lbeas and V. Aleixandre. A two-variables function generator. Mathematics and Computers in Simulation, 2(1):5–12, January 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800031.

DeBacker:1960:EAD

[5] Ir W. De Backer. Error analysis of a d-c integrator. Mathematics and Computers in Simulation, 2 (1):13–23, January 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800043.

Anonymous:1960:Ia

[6] Anonymous. Informations. Mathematics and Computers in Simulation, 2(1):24–26, January 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800055.

Anonymous:1960:CBa

[7] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 2(1):27–42, January 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800067.

Anonymous:1960:RAa

[8] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 2(1):43, January 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800079.

Anonymous:1960:AL

[9] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 2(1):44–46, January 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800080.

Anonymous:1960:DRa

[10] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 2(1):47–51, January 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800092.

Anonymous:1960:PA

[11] Anonymous. Pages 54–104 (April 1960). Mathematics and Computers in Simulation, 2(2):??, April 1960. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Burnay:1960:E

[12] R. Burnay. Editorial. Mathematics and Computers in Simulation, 2(2):54–55, April 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800109.

Anonymous:1960:E

[13] Anonymous.  Editorial.  Mathe-

matics and Computers in Simulation, 2(2):56, April 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800110.

Hacques:1960:ARD

[14] Gérard Hacques. Analogie rhéoélectrique en domaine illimité de la fonction associée d’un potentiel harmonique à symétrie axiale. (French) [Rheoelectric analogy in unlimited domain of the associated function of a harmonic potential with axial symmetry]. Mathematics and Computers in Simulation, 2(2):57–62, April 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800122.

Kohler:1960:NLF

[15] Heinz Köhler. Die Netzwerkrechenmaschine LHB: Der Firma Linke-Hofmann-Busch Salzgitter-Watenstedt. (German) [The network computer LHB: the company Linke-Hofmann-Busch Salzgitter-Watenstedt]. Mathematics and Computers in Simulation, 2(2):63–68, April 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800134.

Renard:1960:NAD

[16] Guy Renard. Nouvelles applications du principe d’inversion dans le calcul analogique experimental. Mathematics and Computers in Simulation, 2(2):69–72, April 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800146.

Anonymous:1960:Ib

[17] Anonymous. Informations. Mathematics and Computers in Simulation, 2(2):73–75, April 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800158.

Anonymous:1960:CBb

[18] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 2(2):76–91, April 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546080016X.

Anonymous:1960:RAb

[19] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 2(2):92–93, April 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800171.

Anonymous:1960:AAa

[20] Anonymous. Activité de l’association. Mathematics and Computers in Simulation, 2(2):94–96, April 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800183.

Anonymous:1960:DRb

[21] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 2(2):97–104, April 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800195.

Anonymous:1960:PJb

[22] Anonymous. Pages 106–163 (July 1960). Mathematics and Computers in Simulation, 2(3):??, July 1960. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

McCann:1960:NDA

[23] G. D. McCann. New developments in analog computation. Mathematics and Computers in Simulation, 2 (3):106–107, July 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800201.

McCann:1960:PRC

[24] G. D. McCann and J. H. Trad. Progrès récents en calcul analogique. (French) [recent progress in analog calculation]. Mathematics and Computers in Simulation, 2(3):108–109, July 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800213.

Hinds:1960:ATV

[25] G. H. Hinds. About technical vocabulary. Mathematics and Computers in Simulation, 2(3):110, July 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800225.

Hinds:1960:PDV

[26] G. H. Hinds. A propos du Vocabulaire technique. Mathematics and Computers in Simulation, 2(3):111, July 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800237.

Herrent:1960:EAR

[27] A. Herrent. Etude analogique d’une regulation de densite dans un processus hydrometallurgique. Mathematics and Computers in Simulation, 2(3):112–121, July 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800249.

Petrich:1960:ASD

[28] J. Petrich and P. Madich. Analytical solution of differential equations with the use of differential analyser. Mathematics and Computers in Simulation, 2(3):122–126, July 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800250.

Raymond:1960:MIF

[29] F.-H. Raymond. Methode d’integration fractionnee. Mathematics and Computers in Simulation, 2(3):127–131, July 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475460800262.

Anonymous:1960:Ic

[30] Anonymous. Informations. Mathematics and Computers in Simulation, 2(3):132–137, July 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800274.

Anonymous:1960:CBc

[31] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 2(3):138–150, July 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800286.

Anonymous:1960:RAc

[32] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 2(3):151, July 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800298.

Anonymous:1960:AAb

[33] Anonymous. Activité de l’association. Mathematics and Computers in Simulation, 2(3):152–160, July 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800304.

Anonymous:1960:DRc

[34] Anonymous. Documents Reçus. Mathematics and Computers in Simulation, 2(3):161–163, July 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800316.

Anonymous:1960:PO

[35] Anonymous. Pages 166–227 (October 1960). Mathematics and Computers in Simulation, 2(4):??, October 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Redshaw:1960:ACC

[36] S. C. Redshaw. Analogue computation in civil engineering. Mathematics and Computers in Simulation, 2 (4):166–167, October 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800328.

Trad:1960:CAG

[37] J. H. Trad. Le calcul analogique en genie civil. Mathematics and Computers in Simulation, 2(4):168– 169, October 1960. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546080033X.

Kuntze:1960:VLS

[38] Karlheinz Kuntze. Zur Verfahrenstecnik für die Losung von Schwingungsproblemen Mittels eines elektronischen Analogrechners. (German) [On the process technique for the solution of vibration problems using an electronic analog calculator]. Mathematics and Computers in Simulation, 2 (4):170–180, October 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800341.

Raymond:1960:QRI

[39] François H. Raymond. Quelques remarques sur l’integration du mouvement autour du centre de gravite dans un simulateur. Mathematics and Computers in Simulation, 2(4): 181–189, October 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800353.

Vuckovic:1960:SPS

[40] V. Vuckovic, P. Miljanic, and I. Obradovic. Solving power system problems by means of the a-c-network analyzer with current sources. Mathematics and Computers in Simulation, 2(4): 190–197, October 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800365.

Anonymous:1960:Id

[41] Anonymous. Informations. Mathematics and Computers in Simulation, 2 (4):198–201, October 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800377.

Anonymous:1960:CBd

[42] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 2(4):202–211, October 1960. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800389.

Anonymous:1960:RAd

[43] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 2(4):211–212, October 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475460800390.

Kovach:1960:AL

[44] Ladis D. Kovach. Activité de l’Association. Mathematics and Computers in Simulation, 2(4):212– 215, October 1960. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800407.

Anonymous:1960:DRd

[45] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 2(4):216–224, October 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800419.

Anonymous:1960:IS

[46] Anonymous. Index — Sommaire. Mathematics and Computers in Simulation, 2(4):225–227, October 1960. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475460800420.

Anonymous:1961:PJa

[47] Anonymous. Pages 2–51 (January 1961). Mathematics and Computers in Simulation, 3(1):??, January 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Verhagen:1961:HEA

[48] C. J. D. M. Verhagen. Higher education and analogue computation. Mathematics and Computers in Simulation, 3(1):2–3, January 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800116.

Trad:1961:ESC

[49] J. H. Trad. Enseignement supérieur et calcul analogique. Mathematics and Computers in Simulation, 3 (1):4–5, January 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800128.

Green:1961:PSB

[50] Francis P. Green. Pacit, a small-budget expandable precision analog computer. Mathematics and Computers in Simulation, 3(1):6–10, January 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546180013X.

Petrich:1961:TMC

[51] J. Petrich. Trigonometric method for conformal mapping with algebraic polynomials by the use of a repetitive differential analyser. Mathematics and Computers in Simulation, 3 (1):11–17, January 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800141.

Petternella:1961:DGF

[52] M. Petternella and A. Ruberti. A diode generator of functions of two variables. Mathematics and Computers in Simulation, 3(1):18–24, January 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800153.

Anonymous:1961:Ia

[53] Anonymous. Informations. Mathematics and Computers in Simulation, 3(1):25–28, January 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800165.

Anonymous:1961:CBa

[54] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 3(1):29–41, January 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800177.

Anonymous:1961:RAa

[55] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 3(1):42, January 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800189.

Anonymous:1961:DRa

[56] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 3(1):43–48, January 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800190.

Anonymous:1961:ALa

[57] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 3(1):49–51, January 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800207.

Anonymous:1961:PA

[58] Anonymous. Pages 52–104 (April 1961). Mathematics and Computers in Simulation, 3(2):??, April 1961. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Braffort:1961:CAD

[59] P. Braffort. Le calcul analogique devant les besoins du genie atomique. Mathematics and Computers in Simulation, 3(2):52–53, April 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800219.

Braffort:1961:ACA

[60] P. Braffort. Analogue computation applied to the needs of atomic engineering. Mathematics and Computers in Simulation, 3(2):54–55, April 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800220.

Hennig:1961:CTW

[61] Tore Hennig. Comparison of two ways of simulating a simple tubular heat exchanger. Mathematics and Computers in Simulation, 3(2):56–59, April 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800232.

Ach:1961:ACT

[62] Ladis D. Kov Ach and William Comley. The analog computer as a teaching aid in diffential equations. Mathematics and Computers in Simulation, 3(2):60–63, April 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800244.

Savastano:1961:SSR

[63] G. Savastano. The simulation of saturable reactors with digital differential analyzers. Mathematics and Computers in Simulation, 3(2): 64–73, April 1961. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800256.

Anonymous:1961:Ib

[64] Anonymous. Informations. Mathematics and Computers in Simulation, 3(2):74–76, April 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800268.

Anonymous:1961:CBb

[65] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 3(2):77–92, April 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546180027X.

Anonymous:1961:RAb

[66] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 3(2):93, April 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475461800281.

Anonymous:1961:DRb

[67] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 3(2):94–99, April 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475461800293.

Anonymous:1961:ALb

[68] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 3(2):100–104, April 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546180030X.

Anonymous:1961:PJb

[69] Anonymous. Pages 106–159 (July 1961). Mathematics and Computers in Simulation, 3(3):??, July 1961. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Coales:1961:ACA

[70] J. F. Coales. Analogue computing and automatic control. Mathematics and Computers in Simulation, 3 (3):106–107, July 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475461800013.

Hoffmann:1961:CAR

[71] Trad. J. Hoffmann. Calcul analogique et réglage automatique. Mathematics and Computers in Simulation, 3 (3):108–109, July 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800025.

Earls:1961:BSD

[72] J. C. Earls. Backlash simulation in D.C. electronic analogue computers. Mathematics and Computers in Simulation, 3(3):110–113, July 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800037.

Hannakam:1961:EDB

[73] L. Hannakam. Ermittlung des dynamischen Betriebsverhaltens elektrischer Maschinen mit Hilfe des elektronischen Analogrechners. (German) [Determination of the dynamic operating behavior of electrical machines using the electronic analog calculator]. Mathematics and Computers in Simulation, 3(3):114–123, July 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800049.

Solbakken:1961:TCF

[74] A. Solbakken. A transistorized carrier frequency computing amplifier. Mathematics and Computers in Simulation, 3(3):124–131, July 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475461800050.

Anonymous:1961:Ic

[75] Anonymous. Informations. Mathematics and Computers in Simulation, 3(3):132–135, July 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475461800062.

Anonymous:1961:CBc

[76] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 3(3):136–149, July 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475461800074.

Anonymous:1961:RAc

[77] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 3(3):150, July 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475461800086.

Anonymous:1961:DRc

[78] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 3(3):151–156, July 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475461800098.

Anonymous:1961:ALc

[79] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 3(3):157–159, July 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800104.

Anonymous:1961:PO

[80] Anonymous. Pages 162–208 (October 1961). Mathematics and Computers in Simulation, 3(4):??, October 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tomovic:1961:E

[81] Rajko Tomovic. Editorial. Mathematics and Computers in Simulation, 3(4):162, October 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800311.

Hoffmann:1961:E

[82] Trad. J. Hoffmann. Editorial. Mathematics and Computers in Simulation, 3(4):163, October 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800323.

Goldenberg:1961:DEF

[83] H. Goldenberg. The determination of electromagnetic forces in transformer windings with axial symmetry by means of a resistance network. Mathematics and Computers in Simulation, 3 (4):164–166, October 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800335.

Hartmann:1961:LCT

[84] John Hartmann, Granino A. Korn, and Richard L. Maybach. Lowcost triangle-integration multipliers for analog computers. Mathematics and Computers in Simulation, 3 (4):167–172, October 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800347.

Hannakam:1961:NES

[85] L. Hannakam. Nachbildung der elektromechanischen Systemgleichungen auf dem elektronischen Analogrechner. (German) [Replica of the electromechanical system equations on the electronic analog calculator]. Mathematics and Computers in Simulation, 3 (4):173–187, October 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800359.

Anonymous:1961:Id

[86] Anonymous. Informations. Mathematics and Computers in Simulation, 3 (4):188–190, October 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800360.

Anonymous:1961:CBd

[87] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 3(4):191–199, October 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800372.

Anonymous:1961:RAd

[88] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 3(4):200, October 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800384.

Anonymous:1961:DRd

[89] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 3(4):201–204, October 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800396.

Anonymous:1961:ALd

[90] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 3(4):205–207, October 1961. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800402.

Anonymous:1961:E

[91] Anonymous.  Errata.  Mathemat-

ics and Computers in Simulation, 3 (4):207, October 1961. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475461800414.

Anonymous:1962:PJa

[92] Anonymous. Pages 2–72 (January 1962). Mathematics and Computers in Simulation, 4(1):??, January 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Korn:1962:ACL

[93] Granino A. Korn. The analog-computer laboratory and the future. Mathematics and Computers in Simulation, 4(1):2–3, January 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800123.

Traduction:1962:LDC

[94] J. H. Traduction. Le laboratoire du calcul analogique et l’avenir. Mathematics and Computers in Simulation, 4(1):4–5, January 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800135.

Witsenhausen:1962:RAE

[95] H. Witsenhausen. Representation analogique des échanges thermiques au sein de réacteurs nucléaires hétérogènes. (French) [Analog representation of thermal exchanges within heterogeneous nuclear reactors]. Mathematics and Computers in Simulation, 4 (1):6–14, January 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800147.

Butler:1962:DAT

[96] R. Butler, E. C. Lloyd, J. G. L. Michel, and Elizabeth D. Sully. Differential analyser treatment of the problem of solidifying steel. Mathematics and Computers in Simulation, 4 (1):15–26, January 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800159.

Kulisch:1962:BDI

[97] Ulrich Kulisch. Behandlung von differentialgleichungen im komplexen auf dem elektronischen analogrechner. (German) [Treatment of differential equations in complex systems on electronic analog calculators]. Mathematics and Computers in Simulation, 4 (1):27–37, January 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800160.

Adt:1962:SMA

[98] Lucien W. Neust Adt. Simulating money allocation problems on an analog computer. Mathematics and Computers in Simulation, 4 (1):38–39, January 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800172.

Anonymous:1962:Ia

[99] Anonymous. Informations. Mathematics and Computers in Simulation, 4(1):40–44, January 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800184.

Anonymous:1962:CBa

[100] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 4(1):45–53, January 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800196.

Anonymous:1962:RAa

[101] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 4(1):54–55, January 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800202.

Anonymous:1962:DRa

[102] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 4(1):56–63, January 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800214.

Anonymous:1962:ALa

[103] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 4(1):64–72, January 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800226.

Anonymous:1962:PA

[104] Anonymous. Pages 74–130 (April 1962). Mathematics and Computers in Simulation, 4(2):??, April 1962. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

DeWinne:1962:CAE

[105] Arthur L. De Winne. Le calcul analogique et l’enseignement des mathematiques appliquees. Mathematics and Computers in Simulation, 4(2):74–75, April 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800354.

DeWinne:1962:ACA

[106] Arthur L. De Winne. Analogue computation and applied mathematics training. Mathematics and Computers in Simulation, 4(2):76–77, April 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800366.

Gilliland:1962:MCF

[107] M. C. Gilliland. A mutually coupled finite difference operator for diffusion processes. Mathematics and Computers in Simulation, 4(2):78–82, April 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800378.

Ja:1962:SSO

[108] Kogan B. Ja. On the steady-state operational amplifier errors due to the pass band limitations. Mathematics and Computers in Simulation, 4(2):83–85, April 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546280038X.

Prawitz:1962:SAT

[109] Jan Prawitz. Simulation of an atmospheric transport problem. Mathematics and Computers in Simulation, 4(2):86–92, April 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800391.

DeWinne:1962:UMC

[110] Arthur L. De Winne. On the use of memory circuits in analogue computers. Mathematics and Computers in Simulation, 4(2):93–105, April 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800408.

Anonymous:1962:TPO

[111] Anonymous. Three phase oscillator. Mathematics and Computers in Simulation, 4(2):106, April 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546280041X.

Anonymous:1962:GAF

[112] Anonymous. Generation of arbitrary functions of time. Mathematics and Computers in Simulation, 4(2):107, April 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800421.

Anonymous:1962:Ib

[113] Anonymous. Informations. Mathematics and Computers in Simulation, 4(2):108–111, April 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800433.

Anonymous:1962:CBb

[114] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 4(2):112–119, April 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800445.

Anonymous:1962:RAb

[115] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 4(2):120–121, April 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800457.

Anonymous:1962:ALb

[116] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 4(2):121–125, April 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800469.

Anonymous:1962:DRb

[117] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 4(2):126–130, April 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800470.

Anonymous:1962:PJb

[118] Anonymous. Pages 132–174 (July 1962).

Mathematics and Computers in Simulation, 4(3):??, July 1962. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Herrmann:1962:AKR

[119] Horst Herrmann. Automatische Kommunikation mit Rechenanlagen. (German) [Automatic communication with computers]. Mathematics and Computers in Simulation, 4(3):132– 133, July 1962. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800019.

Wadhwa:1962:PDO

[120] L. K. Wadhwa. Phase distortion in operational amplifiers. Mathematics and Computers in Simulation, 4 (3):134–137, July 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800020.

Fuchs:1962:SST

[121] A. Fuchs and H. Gafni. A solid state time-division multiplier. Mathematics and Computers in Simulation, 4(3):138–142, July 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800032.

Martin:1962:MON

[122] J. W. Martin, E. F. Harding, and J. Saunders. A method of obtaining noise sources of variable correlation on the pace analogue computer. Mathematics and Computers in Simulation, 4(3):143–148, July 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800044.

Witsenhausen:1962:CDS

[123] Hans Witsenhausen. Circuit for discrete summation. Mathematics and Computers in Simulation, 4(3):149– 150, July 1962. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800056.

Witsenhausen:1962:SSF

[124] Hans Witsenhausen. Shock situation in Fourier-type equations. Mathematics and Computers in Simulation, 4(3):150–151, July 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800068.

Anonymous:1962:Ic

[125] Anonymous. Informations. Mathematics and Computers in Simulation, 4(3):152–156, July 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546280007X.

Anonymous:1962:CBc

[126] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 4(3):157–165, July 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800081.

Anonymous:1962:RAc

[127] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 4(3):165, July 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800093.

Anonymous:1962:ALc

[128] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 4(3):166–168, July 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546280010X.

Anonymous:1962:DRc

[129] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 4(3):170–174, July 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800111.

Anonymous:1962:PO

[130] Anonymous. Pages 176–233 (October 1962). Mathematics and Computers in Simulation, 4(4):??, October 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Denis:1962:E

[131] F. Denis.  Editorial.  Mathemat-

ics and Computers in Simulation, 4 (4):176–178, October 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800238.

Rasof:1962:ACA

[132] B. Rasof and R. Tomovic. The analog computer as an aid in teaching mathematics. Mathematics and Computers in Simulation, 4(4):179– 183, October 1962. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546280024X.

Peretz:1962:FNL

[133] Richard Peretz. Fonctions non linéaires à plusieurs variables. (French) [nonlinear functions with several variables]. Mathematics and Computers in Simulation, 4(4):184–196, October 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800251.

Korn:1962:ALS

[134] Granino A. Korn. Adjoint linear systems in analog computation: a new look. Mathematics and Computers in Simulation, 4(4):197–201, October 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800263.

Vichnevetsky:1962:ETL

[135] R. Vichnevetsky. An extension of a theorem of linear circuit theory to discontinuous delay simulators. Mathematics and Computers in Simulation, 4(4):202, October 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800275.

Earls:1962:GLF

[136] J. C. Earls. The generation of lowfre quency triangular voltages. Mathematics and Computers in Simulation, 4(4):203–204, October 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800287.

Laws:1962:I

[137] R. A. Laws. Informations. Mathematics and Computers in Simulation, 4(4):205–207, October 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800299.

Anonymous:1962:CBd

[138] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 4(4):208–218, October 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800305.

Anonymous:1962:RAd

[139] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 4(4):219–220, October 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800317.

Anonymous:1962:ALd

[140] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 4(4):221–225, October 1962. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475462800329.

Anonymous:1962:DRd

[141] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 4(4):226–231, October 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800330.

Anonymous:1962:IS

[142] Anonymous. Index — Sommaire. Mathematics and Computers in Simulation, 4(4):232–233, October 1962. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475462800342.

Anonymous:1963:PJa

[143] Anonymous. Pages 2–57 (January 1963).

Mathematics and Computers in Simulation, 5(1):??, January 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Karplus:1963:TFS

[144] Walter J. Karplus. Trends in field simulation. Mathematics and Computers in Simulation, 5(1):2–3, January 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800270.

Paschkis:1963:HMF

[145] V. Paschkis. The heat and mass flow analyzer a simulator for the study of heat conduction. Mathematics and Computers in Simulation, 5(1):4–18, January 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800282.

Vidal:1963:RIT

[146] Jacques J. Vidal and Gilbert M. H. Burnay. Les reseaux d’impedances Tendances et developpements recents. Mathematics and Computers in Simulation, 5(1):19–26, January 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800294.

Vidal:1963:AEE

[147] Jacques J. Vidal. Application des effets élémentaires à la solution analogiquedigitale d’un problème stationnaire à condition aux limites non linéaire. (French) [application of elementary effects to the analog-digital solution of a stationary problem provided with nonlinear limits]. Mathematics and Computers in Simulation, 5 (1):27–28, January 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800300.

Anonymous:1963:Ia

[148] Anonymous. Informations. Mathematics and Computers in Simulation, 5(1):29–30, January 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800312.

Anonymous:1963:CBa

[149] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 5(1):31–37, January 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800324.

Anonymous:1963:RAa

[150] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 5(1):38, January 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800336.

Anonymous:1963:ALa

[151] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 5(1):39–44, January 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800348.

Anonymous:1963:DRa

[152] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 5(1):45–57, January 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546380035X.

Anonymous:1963:PA

[153] Anonymous. Pages 60–130 (April 1963).

Mathematics and Computers in Simulation, 5(2):??, April 1963. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Gleitz:1963:E

[154] J. J. Gleitz.  Editorial.  Math-

ematics and Computers in Simulation, 5(2):60, April 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800117.

Millar:1963:E

[155] J. F. Millar.  Editorial.  Math-

ematics and Computers in Simulation, 5(2):61, April 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800129.

Karplus:1963:LE

[156] Walter J. Karplus. Letter to the editor. Mathematics and Computers in Simulation, 5(2):62, April 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800130.

Green:1963:CAI

[157] C. Green, A. Debroux, G. P. Del Bigio, and H. D’Hoop. The code Apache intended for the programming of an analogue problem by means of a digital computer. Mathematics and Computers in Simulation, 5(2):63–74, April 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800142.

Noda:1963:HCA

[158] Katsuhiko Noda, Saburo Tsuji, and Noboru Sugie. The hybrid computer for automatic control. Mathematics and Computers in Simulation, 5(2):75–79, April 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800154.

Vichnevetsky:1963:DPH

[159] R. Vichnevetsky and J. P. Waha. Dynamic programming and hybrid computation. Mathematics and Computers in Simulation, 5(2):80–87, April 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800166.

Vichnevetsky:1963:ECB

[160] R. Vichnevetsky. Experimental calibration of band pass filters by the impulse method. Mathematics and Computers in Simulation, 5(2):88–89, April 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800178.

Korn:1963:AMC

[161] Granino A. Korn. Analog multicomparator for multiple decisions or analog-to-digital conversion. Mathematics and Computers in Simulation, 5(2):89–90, April 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546380018X.

Bayly:1963:UBE

[162] J. G. Bayly and A. C. Soudack.

Use of the Berkeley Easecomputer in evaluating double integrals. Mathematics and Computers in Simulation, 5(2):90–93, April 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800191.

Korn:1963:PPG

[163] Granino A. Korn. Parameterperturbation generator for analogcomputer error and optimization studies. Mathematics and Computers in Simulation, 5(2):93–95, April 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800208.

Giloi:1963:SLG

[164] W. Giloi. Über Stabilitätsschwierigkeiten bei der Lösung Gewisser Fälle der Diffusions-und Wärmeleitungsgleichung Auf Dem Analogrechner. (German) [On stabilizer difficulties in the solution of certain cases of the diffusion and heat conduction equation on the analog calculator]. Mathematics and Computers in Simulation, 5(2):95–98, April 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546380021X.

Anonymous:1963:Ib

[165] Anonymous. Informations. Mathematics and Computers in Simulation, 5(2):99–105, April 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800221.

Anonymous:1963:CBb

[166] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 5(2):106–113, April 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800233.

Anonymous:1963:RAb

[167] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 5(2):114–115, April 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800245.

Anonymous:1963:ALb

[168] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 5(2):116–120, April 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800257.

Anonymous:1963:DRb

[169] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 5(2):121–130, April 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800269.

Anonymous:1963:PJb

[170] Anonymous. Pages 132–192 (July 1963).

Mathematics and Computers in Simulation, 5(3):??, July 1963. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vichnevetsky:1963:PCP

[171] R. Vichnevetsky. Pour un contact plus étroit entre les professionnels du calcul analogique. (French) [for more and more contact between professionals in analog calculation]. Mathematics and Computers in Simulation, 5(3):132–133, July 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800361.

Millar:1963:TCC

[172] J. F. Millar. Towards a closer contact between analogue Computation specialists. Mathematics and Computers in Simulation, 5(3):134–135, July 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800373.

Anonymous:1963:LE

[173] Anonymous. Letter to the editor. Mathematics and Computers in Simulation, 5(3):136, July 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800385.

Petric:1963:ARD

[174] Jovan Petric. The application of a repetitive differential analyzer to a flutter problem. Mathematics and Computers in Simulation, 5(3):137– 140, July 1963. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800397.

Miura:1963:TSC

[175] Takeo Miura and Junzo Iwata. Timesharing computation utilizing analog memory. Mathematics and Computers in Simulation, 5(3):141– 149, July 1963. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800403.

Sylin:1963:TAT

[176] G. Sylin. Traceur automatique de trajectoires de particules chargées. (French) [automatic tracer of trajectories of charged particles]. Mathematics and Computers in Simulation, 5(3):150–155, July 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800415.

Fife:1963:SSR

[177] Dennis W. Fife. Simrar: Simulated receiver and recorder for statistical measurements. Mathematics and Computers in Simulation, 5 (3):156–159, July 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800427.

vanderGrinten:1963:PAC

[178] P. M. E. M. van der Grinten and W. Krijger. Processing of the auto and crosscorrelation functions to step response. Mathematics and Computers in Simulation, 5(3):160–161, July 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800439.

Anonymous:1963:Ic

[179] Anonymous. Informations. Mathematics and Computers in Simulation, 5(3):162–165, July 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475463800440.

Anonymous:1963:CBc

[180] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 5(3):166–178, July 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475463800452.

Anonymous:1963:RAc

[181] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 5(3):179–180, July 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475463800464.

Anonymous:1963:ALc

[182] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 5(3):181–185, July 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475463800476.

Anonymous:1963:DRc

[183] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 5(3):186–192, July 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800488.

Anonymous:1963:PO

[184] Anonymous. Pages 194–241 (October 1963). Mathematics and Computers in Simulation, 5(4):??, October 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Haus:1963:CAA

[185] Frederic Haus. Le calcul analogique et l’aeronautique. Mathematics and Computers in Simulation, 5(4):194– 195, October 1963. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800014.

Hempel:1963:ESS

[186] Arvid B. Hempel. Exact simulation of a special first order partial differential equation. Mathematics and Computers in Simulation, 5 (4):196–199, October 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800026.

King:1963:CPD

[187] R. E. King and W. A. Brown. A cumulative probability distribution analyser. Mathematics and Computers in Simulation, 5(4):200–203, October 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800038.

Dekker:1963:RMD

[188] Ir. L. Dekker and W. Heins. Resonance method for the determination of the eigenvalues of matrices on an analog computer. Mathematics and Computers in Simulation, 5 (4):204–209, October 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546380004X.

Solheim:1963:OTP

[189] Ole Andreas Solheim. Optimum temperature profiles in tubular reactors. Mathematics and Computers in Simulation, 5(4):210–215, October 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475463800051.

Anonymous:1963:Id

[190] Anonymous. Informations. Mathematics and Computers in Simulation, 5 (4):216–219, October 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475463800063.

Anonymous:1963:CBd

[191] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 5(4):220–228, October 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475463800075.

Anonymous:1963:RAd

[192] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 5(4):229, October 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800087.

Anonymous:1963:DRd

[193] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 5(4):230–238, October 1963. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800099.

Anonymous:1963:AA

[194] Anonymous. Activité de l’association.

Mathematics and Computers in Simulation, 5(4):238–241, October 1963. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475463800105.

Anonymous:1964:PJa

[195] Anonymous. Pages 2–74 (January 1964).

Mathematics and Computers in Simulation, 6(1):??, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Rijlant:1964:E

[196] P. Rijlant. Editorial. Mathematics and Computers in Simulation, 6 (1):2–3, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546480001X.

Petric:1964:AEM

[197] Jovan Petric. Analog electrical model for the use of Mitrovic’s criterion in analysis feedback control systems. Mathematics and Computers in Simulation, 6(1):4–8, January 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800021.

Mathis:1964:FDA

[198] A. Mathis and P. Giordano. Finite differences analog models of physical systems containing flowing fluids: Error analysis and design criteria. Mathematics and Computers in Simulation, 6(1):9–21, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800033.

Eckes:1964:DES

[199] Harry R. Eckes. Digital expansion system for ASTRAC I. Mathematics and Computers in Simulation, 6 (1):22–31, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800045.

Mitchell:1964:HAD

[200] Baker A. Mitchell. A hybrid/analogdigital one-parameter optimizer. Mathematics and Computers in Simulation, 6(1):32–37, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800057. See

comment [226].

Soudack:1964:AAC

[201] A. C. Soudack and J. G. Bayly.

Application of the analog computer to waveform analysis. Mathematics and Computers in Simulation, 6 (1):38–40, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800069.

Tunmore:1964:NAT

[202] B. G. Tunmore. A network analyser technique for non-sinusoidal inputs. Mathematics and Computers in Simulation, 6(1):41–42, January 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800070.

Anonymous:1964:RAa

[203] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 6(1):43–44, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800082.

Anonymous:1964:Ia

[204] Anonymous. Informations. Mathematics and Computers in Simulation, 6(1):45–49, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800094.

Anonymous:1964:CBa

[205] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 6(1):50–60, January 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800100.

Anonymous:1964:DRa

[206] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 6(1):61–70, January 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800112.

Anonymous:1964:ALa

[207] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 6(1):71–74, January 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800124.

Anonymous:1964:PA

[208] Anonymous. Pages 76–153 (April 1964). Mathematics and Computers in Simulation, 6(2):??, April 1964. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Soudack:1964:ACE

[209] Avrum C. Soudack. Analog computers and engineering education. Mathematics and Computers in Simulation, 6(2):76, April 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800136.

Brubaker:1964:SHS

[210] Thomas A. Brubaker. A study of highspeed analog-computer performance (the ASTRAC I system performance). Mathematics and Computers in Simulation, 6(2):77–88, April 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800148.

Holst:1964:IAC

[211] Per Asbjörn Holst. Iterative analog computations. Mathematics and Computers in Simulation, 6(2):89– 104, April 1964. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546480015X.

Vichnevetsky:1964:ACS

[212] R. Vichnevetsky. Analogue computer simulation of a time-dependant delay using the concept of generalised transfer function. Mathematics and Computers in Simulation, 6 (2):105–109, April 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800161.

King:1964:RST

[213] R. E. King. A random sampletime generator with defined statistical properties. Mathematics and Computers in Simulation, 6(2):110– 111, April 1964. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800173.

Witsenhausen:1964:DPH

[214] Hans Witsenhausen. Development of a program for the hybrid simulation of a tubular reactor. Mathematics and Computers in Simulation, 6(2):112–117, April 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800185.

Landauer:1964:SSV

[215] J. Paul Landauer. Simulation of space vehicle with reaction jet control system. Mathematics and Computers in Simulation, 6(2):118–124, April 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800197.

Anonymous:1964:RAb

[216] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 6(2):125–126, April 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800203.

Anonymous:1964:Ib

[217] Anonymous. Informations. Mathematics and Computers in Simulation, 6(2):127–131, April 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800215.

Anonymous:1964:CBb

[218] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 6(2):132–138, April 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800227.

Anonymous:1964:DRb

[219] Anonymous. Documents Reçus. Mathematics and Computers in Simulation, 6(2):139–149, April 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800239.

Anonymous:1964:AA

[220] Anonymous. Activité de l’association.

Mathematics and Computers in Simulation, 6(2):150–153, April 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800240.

Anonymous:1964:PJb

[221] Anonymous. Pages 156–227 (July 1964).

Mathematics and Computers in Simulation, 6(3):??, July 1964. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Hoffmann:1964:E

[222] J. Hoffmann.  Editorial.  Math-

ematics and Computers in Simulation, 6(3):156, July 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800367.

Bingulac:1964:OIC

[223] S. Bingulac and E. Humo. Optimalisation iterative sur calculateur analogique.

Mathematics and Computers in Simulation, 6(3):157–161, July 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800379.

Baudras-Chardigny:1964:CUA

[224] A. Baudras-Chardigny. Le calculateur universel analac. (French) [The universal calculator analac]. Mathematics and Computers in Simulation, 6(3):162–168, July 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800380.

Miura:1964:CCS

[225] Takeo Miura and Junzo Iw Ata. Combined computation with slow-type and repetitive analog computers. Mathematics and Computers in Simulation, 6(3):169–177, July 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800392.

Mitchell:1964:CHA

[226] Baker A. Mitchell. Comment on “A hybrid/analog-digital one-parameter optimizer. Mathematics and Computers in Simulation, 6(3):178, July 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800409. See

[200].

Sams:1964:ASE

[227] Sams. Arrangement and scaling of equations. Mathematics and Computers in Simulation, 6(3):179– 182, July 1964. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800410.

Anonymous:1964:RAc

[228] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 6(3):183, July 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800422.

Anonymous:1964:Ic

[229] Anonymous. Informations. Mathematics and Computers in Simulation, 6(3):184–193, July 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800434.

Anonymous:1964:CBc

[230] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 6(3):194–217, July 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800446.

Anonymous:1964:DRc

[231] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 6(3):218–224, July 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800458.

Anonymous:1964:ALb

[232] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 6(3):225–227, July 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546480046X.

Anonymous:1964:PO

[233] Anonymous. Pages 230–282 (October 1964). Mathematics and Computers in Simulation, 6(4):??, October 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Baleriaux:1964:CAI

[234] H. Baleriaux. Le calcul analogique et l’industrie de la production, du transport et de la distribution de l’energie electrique. Mathematics and Computers in Simulation, 6(4): 230–231, October 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800252.

Noronha:1964:AST

[235] L. G. Noronha. Analogue simulation of transfer functions. Mathematics and Computers in Simulation, 6 (4):232–242, October 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800264.

Wiig:1964:ESD

[236] Karl M. Wiig. An evaluation of static and dynamic errors in the solution of a first-order partial differential equation on an analog computer. Mathematics and Computers in Simulation, 6(4):243–248, October 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800276.

Devooght:1964:MID

[237] J. Devooght and G. Sylin. Mesure de l’intégrale de Dirichlet et ses applications. (French) [Evaluation of the Dirchlet integral and its applications]. Mathematics and Computers in Simulation, 6(4):249–255, October 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800288.

Causon:1964:SRT

[238] A. Causon. Some roulette-type curves.

Mathematics and Computers in Simulation, 6(4):256–258, October 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546480029X.

Anonymous:1964:RAd

[239] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 6(4):259, October 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475464800306.

Anonymous:1964:Id

[240] Anonymous. Informations. Mathematics and Computers in Simulation, 6 (4):260–261, October 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800318.

Anonymous:1964:CBd

[241] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 6(4):262–268, October 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546480032X.

Anonymous:1964:DRd

[242] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 6(4):269–275, October 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800331.

Anonymous:1964:ALc

[243] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 6(4):276–279, October 1964. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800343.

Anonymous:1964:IS

[244] Anonymous. Index — sommaire. Mathematics and Computers in Simulation, 6(4):280–282, October 1964. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475464800355.

Anonymous:1965:PJa

[245] Anonymous. Pages 2–68 (January 1965).

Mathematics and Computers in Simulation, 7(1):??, January 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Honnell:1965:MMT

[246] Pierre H. Honnell. Matric mathematics, theory, and computers. Mathematics and Computers in Simulation, 7(1):2–3, January 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800015.

Visman:1965:DFG

[247] A. C. P. Visman. A diode function generator made suitable for two variables. Mathematics and Computers in Simulation, 7(1):4–6, January 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800027.

Pesina:1965:DCR

[248] I. Pesina. Determining the character of the relation between heat input and reflux ratio in a rectification column. Mathematics and Computers in Simulation, 7(1):7–11, January 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800039.

Bingulac:1965:AOL

[249] S. Bingulac and P. Kokotovic. Automatic optimization of linear feedback control systems on an analog computer.

Mathematics and Computers in Simulation, 7(1):12–17, January 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800040.

Douce:1965:ASC

[250] J. L. Douce and H. Wilson. The automatic synthesis of control systems with constraints. Mathematics and Computers in Simulation, 7 (1):18–22, January 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800052.

Earls:1965:DCE

[251] J. C. Earls. A.D.C. electronic analogue simulation of the subharmonic response of a series ferroresonant circuit. Mathematics and Computers in Simulation, 7(1):23–24, January 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800064.

Soudack:1965:ACS

[252] A. C. Soudack and J. T. Walton. An analog computer study of parasitic effects on tunnel-diode switching speeds. Mathematics and Computers in Simulation, 7(1):25–26, January 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800076.

Anonymous:1965:RAa

[253] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 7(1):27, January 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800088.

Anonymous:1965:Ia

[254] Anonymous. Informations. Mathematics and Computers in Simulation, 7(1):28–32, January 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546580009X.

Anonymous:1965:CBa

[255] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 7(1):33–51, January 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800106.

Anonymous:1965:DRa

[256] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 7(1):52–61, January 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800118.

Anonymous:1965:ALa

[257] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 7(1):62–68, January 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546580012X.

Anonymous:1965:PA

[258] Anonymous. Pages 70–128 (April 1965). Mathematics and Computers in Simulation, 7(2):??, April 1965. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ushakov:1965:E

[259] Valentin B. Ushakov. Editorial. Mathematics and Computers in Simulation, 7(2):70–71, April 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800131.

Berry:1965:UAC

[260] F. G. Berry and A. C. Soudack. The use of the analog computer in teaching engineering systems. Mathematics and Computers in Simulation, 7(2):72–76, April 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800143.

Humo:1965:PSI

[261] Emir Humo. Parameter sweep in iterative analog computer techniques. Mathematics and Computers in Simulation, 7(2):77–81, April 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800155.

Krishnamurthy:1965:MAS

[262] E. V. Krishnamurthy and P. M. Honnell. A matric approach to the stability of solutions of algebraic systems by arithmetic and analytic machines.

Mathematics and Computers in Simulation, 7(2):82–91, April 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800167.

Fadden:1965:DNM

[263] Edward J. Fadden. Determination of the normal modes of a two-point boundary value problem. Mathematics and Computers in Simulation, 7(2):92–96, April 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800179.

Anonymous:1965:RAb

[264] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 7(2):96–97, April 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800180.

Anonymous:1965:Ib

[265] Anonymous. Informations. Mathematics and Computers in Simulation, 7(2):98–103, April 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800192.

Anonymous:1965:CBb

[266] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 7(2):104–113, April 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800209.

Anonymous:1965:DRb

[267] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 7(2):114–121, April 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800210.

Anonymous:1965:ALb

[268] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 7(2):122–128, April 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800222.

Anonymous:1965:PJb

[269] Anonymous. Pages 130–174 (July 1965). Mathematics and Computers in Simulation, 7(3):??, July 1965. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Naslin:1965:QCQ

[270] P. Naslin. Qu’est-ce que le calcul analogique? Mathematics and Computers in Simulation, 7(3):130– 131, July 1965. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800234.

Conant:1965:ASO

[271] Brian K. Conant. ASTRAC-I Study of an orthogonal-function communications system. Mathematics and Computers in Simulation, 7(3):132– 137, July 1965. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800246.

Anonymous:1965:RAc

[272] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 7(3):137, July 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800258.

Sevely:1965:CEC

[273] Y. Sevely and P. Vidal. Contribution a l’etude sur calculatrice analogique des asservissements a redresseurs controles et a modulation de largeur d’impulsions. Mathematics and Computers in Simulation, 7(3):138–142, July 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546580026X.

DeGreef:1965:CAE

[274] F. De Greef. Changement automatique d’echelle en calcul analogique. Mathematics and Computers in Simulation, 7(3):143–146, July 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800271.

Anonymous:1965:Ic

[275] Anonymous. Informations. Mathematics and Computers in Simulation, 7(3):147–153, July 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800283.

Anonymous:1965:CBc

[276] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 7(3):154–161, July 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800295.

Anonymous:1965:DRc

[277] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 7(3):162–172, July 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800301.

Anonymous:1965:ALc

[278] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 7(3):173–174, July 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800313.

Anonymous:1965:PO

[279] Anonymous. Pages 176–229 (October 1965). Mathematics and Computers in Simulation, 7(4):??, October 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kokinopoulos:1965:UCE

[280] F. E. Kokinopoulos. Utilisation des calculateurs électroniques analogiques à la recherche antiseismique. (French) [Use of electronic analog computers in antiseismic research]. Mathematics and Computers in Simulation, 7 (4):176, October 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800325.

Boscher:1965:RPR

[281] J. Boscher. Resolution par reseaux resistifs de l’equation de la diffusion. Mathematics and Computers in Simulation, 7(4):177–185, October 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800337.

Dekker:1965:MIH

[282] L. Dekker. Multiplying and integrating in a hybrid way. Mathematics and Computers in Simulation, 7 (4):186–189, October 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800349.

Little:1965:ACS

[283] W. D. Little and A. C. Soudack. On the analog computer solution of first-order partial differential equations. Mathematics and Computers in Simulation, 7(4):190–194, October 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800350.

Delchamber:1965:ESE

[284] M. Delchamber. Etude de la stabilite d’ecoulements entre plans paralleles en presence d’un champ magnetique a l’aide d’un calculateur analogique. Mathematics and Computers in Simulation, 7(4):195–204, October 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800362.

Anonymous:1965:RAd

[285] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 7(4):205, October 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800374.

Anonymous:1965:CBd

[286] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 7(4):206–214, October 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800386.

Anonymous:1965:AA

[287] Anonymous. Activité de l’association. Mathematics and Computers in Simulation, 7(4):215–217, October 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800398.

Anonymous:1965:DRd

[288] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 7(4):218–223, October 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475465800404.

Anonymous:1965:Id

[289] Anonymous. Informations. Mathematics and Computers in Simulation, 7 (4):224–229, October 1965. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800416.

Anonymous:1965:RAT

[290] Anonymous. Recommendations to authors of texts for publication in the annals of the international association for analogue computation. Mathematics and Computers in Simulation, 7(4):229, October 1965. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475465800428.

Anonymous:1966:PJa

[291] Anonymous. Pages 2–63 (January 1966).

Mathematics and Computers in Simulation, 8(1):??, January 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sevely:1966:E

[292] Y. Sevely. Editorial. Mathematics and Computers in Simulation, 8 (1):2–3, January 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800423.

VanderVorst:1966:RAC

[293] André Vander Vorst. Representation analogique de la commutation de noyaux magnetiques. Mathematics and Computers in Simulation, 8 (1):4–12, January 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800435.

OGrady:1966:HCD

[294] Emmett Pearse O’Grady. A hybridcode differential analyzer. Mathematics and Computers in Simulation, 8 (1):13–21, January 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800447.

Anonymous:1966:RAa

[295] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 8(1):22, January 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800459.

Anonymous:1966:CBa

[296] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 8(1):23–42, January 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800460.

Anonymous:1966:DRa

[297] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 8(1):43–54, January 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800472.

Anonymous:1966:Ia

[298] Anonymous. Informations. Mathematics and Computers in Simulation, 8(1):54–61, January 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800484.

Anonymous:1966:ALa

[299] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 8(1):62–63, January 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800496.

Anonymous:1966:FIA

[300] Anonymous. Fourth International Analogue Computation Meetings Brighton, 14–18 September 1964. Mathematics and Computers in Simulation, 8 (1):iv, January 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800502.

Anonymous:1966:PA

[301] Anonymous. Pages 66–120 (April 1966).

Mathematics and Computers in Simulation, 8(2):??, April 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Jespers:1966:Ea

[302] P. Jespers. Editorial. Mathematics and Computers in Simulation, 8 (2):66, April 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800320.

Jespers:1966:Eb

[303] P. Jespers. Editorial. Mathematics and Computers in Simulation, 8 (2):67, April 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800332.

Honnell:1966:TFM

[304] Pierre M. Honnell and Robert J. Mulholland. Transfer functions and the matric computor. Mathematics and Computers in Simulation, 8(2):68–77, April 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800344.

Mukherjee:1966:SSS

[305] S. Mukherjee. Study of slow surface wave by analog computor. Mathematics and Computers in Simulation, 8(2):78–84, April 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800356.

Melton:1966:ACS

[306] R. D. Melton and A. C. Soudack. Analog computer study of a non-linear automobile suspension system. Mathematics and Computers in Simulation, 8(2):85–88, April 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800368.

Anonymous:1966:RAb

[307] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 8(2):89, April 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546680037X.

Anonymous:1966:CBb

[308] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 8(2):90–99, April 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800381.

Anonymous:1966:DRb

[309] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 8(2):100–109, April 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800393.

Anonymous:1966:Ib

[310] Anonymous. Informations. Mathematics and Computers in Simulation, 8(2):110–116, April 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546680040X.

Anonymous:1966:ALb

[311] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 8(2):117–120, April 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800411.

Anonymous:1966:PJb

[312] Anonymous. Pages 122–181 (July 1966). Mathematics and Computers in Simulation, 8(3):??, July 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Santesmases:1966:E

[313] J. Garcia Santesmases. Editorial. Mathematics and Computers in Simulation, 8(3):122–123, July 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800137.

Mathis:1966:IMA

[314] A. Mathis, P. Giordano, and G. Gatto. Iterative methods for the analog simulation of physical systems containing flowing fluids. Mathematics and Computers in Simulation, 8 (3):124–137, July 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800149.

Keymeulen:1966:ERI

[315] R. Keymeulen. Étude et réalisation d’un interrupteur à transistors pour calculateur analogique. (French) [Study and construction of a transistors switch for an analog calculator]. Mathematics and Computers in Simulation, 8(3):138–145, July 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800150.

Dekker:1966:ARE

[316] L. Dekker. Adjusting resistors in an electronic way. Mathematics and Computers in Simulation, 8(3): 146–154, July 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800162.

Anonymous:1966:RAc

[317] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 8(3):155–156, July 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800174.

Anonymous:1966:CBc

[318] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 8(3):157–163, July 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800186.

Anonymous:1966:DRc

[319] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 8(3):164–172, July 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800198.

Anonymous:1966:Ic

[320] Anonymous. Informations. Mathematics and Computers in Simulation, 8(3):173–178, July 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800204.

Anonymous:1966:ALc

[321] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 8(3):180–181, July 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800216.

Anonymous:1966:PO

[322] Anonymous. Pages 184–232 (October 1966). Mathematics and Computers in Simulation, 8(4):??, October 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Renard:1966:RCH

[323] G. Renard. Reflexions sur le calcul hybride. Mathematics and Computers in Simulation, 8(4):184– 186, October 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800010.

Lockhart:1966:SSV

[324] G. B. Lockhart and J. C. Earls. A sequentially-switched variable-gain operational amplifier. Mathematics and Computers in Simulation, 8(4): 187–188, October 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800022.

Aubrun:1966:GFE

[325] Michel Aubrun and Jean-Marie Kauffmann. Generateur de fonction exponentielle. Mathematics and Computers in Simulation, 8(4):189– 192, October 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800034.

Maybach:1966:GIF

[326] R. L. Maybach. Generation of inverse functions by the method of steepest descent. Mathematics and Computers in Simulation, 8(4):193– 196, October 1966. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800046.

Lovas:1966:NPT

[327] Laszlo T. Lovas and Avrum C. Soudack.

A novel paper-tape function generator and multiplier. Mathematics and Computers in Simulation, 8(4): 197–202, October 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800058.

Bozicevic:1966:PRN

[328] Juraj Bozicevic and Ratko Gospodnetic. Preliminary report on a new kind of curve follower. Mathematics and Computers in Simulation, 8 (4):203–205, October 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546680006X.

Anonymous:1966:RAd

[329] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 8(4):206–207, October 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475466800071.

Anonymous:1966:CBd

[330] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 8(4):208–219, October 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475466800083.

Anonymous:1966:DRd

[331] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 8(4):220–223, October 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475466800095.

Anonymous:1966:Id

[332] Anonymous. Informations. Mathematics and Computers in Simulation, 8 (4):224–227, October 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475466800101.

Anonymous:1966:AA

[333] Anonymous. Activité de l’association. Mathematics and Computers in Simulation, 8(4):228–230, October 1966. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800113.

Anonymous:1966:ISV

[334] Anonymous. Index — Sommaire volume VII (1965). Mathematics and Computers in Simulation, 8(4): 231–232, October 1966. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475466800125.

Anonymous:1967:PJa

[335] Anonymous. Pages 2–63 (January 1967).

Mathematics and Computers in Simulation, 9(1):??, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vidal:1967:E

[336] Pierre Vidal.  Editorial.  Math-

ematics and Computers in Simulation, 9(1):2, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800247.

Mizzi:1967:ECA

[337] F. Mizzi and P. Maumet. Étude sur calculateur analogique d’un champ laplacien plan à donneés de Dirichlet: écoulement autour d’un profil variable. (French) [Study on an analog calculator of a planar Laplacian field with Dirichlet data: flow around a variable profile]. Mathematics and Computers in Simulation, 9(1):3–13, January 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800259.

Weidner:1967:ESA

[338] Helmut Weidner. Ein elektronischer Schalter für Analogrechner. (German) [An electronic switch for analog calculator]. Mathematics and Computers in Simulation, 9(1):14– 15, January 1967. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800260.

Paker:1967:APD

[339] Yakup Paker and Stephen H. Unger. ADAC — a programmed direct analog computer. Mathematics and Computers in Simulation, 9(1):16– 23, January 1967. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800272.

Freeman:1967:ACS

[340] E. A. Freeman and M. R. Karaksi. The analogue computer simulation of an alloy junction transistor wide band equivalent circuit. Mathematics and Computers in Simulation, 9 (1):24–29, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800284.

McCann:1967:STD

[341] Michael J. McCann. Simulation techniques for distributed parameter systems. Mathematics and Computers in Simulation, 9(1):30– 33, January 1967. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800296.

Meissinger:1967:IEA

[342] H. F. Meissinger. Interrelation of error analysis, sensitivity analysis, and parameter identification. Mathematics and Computers in Simulation, 9(1):34, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800302.

Korn:1967:ACE

[343] Granino A. Korn. Automated computing-error studies in fast hybrid computation. Mathematics and Computers in Simulation, 9(1): 34, January 1967. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800314.

Karplus:1967:EAH

[344] Walter J. Karplus. Error analysis of hybrid solutions of partial differential equations. Mathematics and Computers in Simulation, 9 (1):34–35, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800326.

Gilbert:1967:EAH

[345] Elmer G. Gilbert. Error-analysis of hybrid computer loops. Mathematics and Computers in Simulation, 9 (1):35–36, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800338.

Illiland:1967:SEA

[346] M. C. Illiland. Spectral error analysis of hybrid subroutines. Mathematics and Computers in Simulation, 9(1):36, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546780034X.

Bekey:1967:AED

[347] George A. Bekey. The analysis of errors due to sampling rate variation. Mathematics and Computers in Simulation, 9(1):36, January 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800351.

Anonymous:1967:RAa

[348] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 9(1):37–38, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800363.

Anonymous:1967:CBa

[349] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 9(1):39–45, January 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800375.

Anonymous:1967:DRa

[350] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 9(1):46–53, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800387.

Anonymous:1967:Ia

[351] Anonymous. Informations. Mathematics and Computers in Simulation, 9(1):54–62, January 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800399.

Anonymous:1967:AA

[352] Anonymous. Activité de l’association.

Mathematics and Computers in Simulation, 9(1):63, January 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800405.

Anonymous:1967:PA

[353] Anonymous. Pages 66–112 (April 1967).

Mathematics and Computers in Simulation, 9(2):??, April 1967. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Gyergyek:1967:CDD

[354] Ludvik Gyergyek. Le choix depend du but et du critere. Mathematics and Computers in Simulation, 9(2):66–67, April 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800417.

Borsei:1967:ACS

[355] A. Borsei and G. Estrin. An analog computer study of the dynamic behavior of stressed thin ferromagnetic films. Mathematics and Computers in Simulation, 9(2):68–75, April 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800429.

Santesmases:1967:ASF

[356] J. G. Santesmases, J. Ayala, and A. H. Cachero. Analog simulation of a ferroresonant system including analysis of hysteresis loop. Mathematics and Computers in Simulation, 9(2):76–80, April 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800430.

Crosbie:1967:ACS

[357] R. E. Crosbie and N. G. Hingorani. Analog computer solution of ship’s deviation due to submarine d.c. cable. Mathematics and Computers in Simulation, 9(2):81–85, April 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800442.

Anonymous:1967:RAb

[358] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 9(2):86–87, April 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800454.

Anonymous:1967:CBb

[359] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 9(2):88–96, April 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800466.

Anonymous:1967:DRb

[360] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 9(2):97–104, April 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800478.

Anonymous:1967:Ib

[361] Anonymous. Informations. Mathematics and Computers in Simulation, 9(2):105–110, April 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546780048X.

Anonymous:1967:ALa

[362] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 9(2):111–112, April 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800491.

Anonymous:1967:PJb

[363] Anonymous. Pages 114–170 (July 1967).

Mathematics and Computers in Simulation, 9(3):??, July 1967. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Cuenod:1967:VCL

[364] M. Cuénod. Le V ^{e} Congrès de l’AICA. Mathematics and Computers in Simulation, 9(3):114–117, July 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800016.

Bajenesco:1967:NRO

[365] T. I. Bajenesco. Nouveaux resultats obtenus a l’aide du calculateur analogique dans l’etude de l’equation du mouvement des particules d’une suspension solide entrainees par un flux courbe. Mathematics and Computers in Simulation, 9(3):118–121, July 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800028.

P:1967:MMI

[366] E. S. I. P. A mathematical model for the investigation of three-dimensional fields with asymmetric boundaries. Mathematics and Computers in Simulation, 9(3):122–130, July 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546780003X.

Kokinopoulos:1967:DADa

[367] E. H. Kokinopoulos. Dynamic aseismic design of an elevated water tower. Mathematics and Computers in Simulation, 9(3):131–141, July 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800041.

Vansteenkiste:1967:OGA

[368] Ch. J. Vansteenkiste. Optimal generation of arbitrary functions. Mathematics and Computers in Simulation, 9(3):142–146, July 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800053.

Anonymous:1967:RAc

[369] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 9(3):147–148, July 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800065.

Anonymous:1967:CBc

[370] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 9(3):149–160, July 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800077.

Anonymous:1967:DRc

[371] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 9(3):161–163, July 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800089.

Anonymous:1967:Ic

[372] Anonymous. Informations. Mathematics and Computers in Simulation, 9(3):164–168, July 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800090.

Anonymous:1967:ALb

[373] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 9(3):169–170, July 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800107.

Anonymous:1967:PO

[374] Anonymous. Pages 172–223 (October 1967). Mathematics and Computers in Simulation, 9(4):??, October 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Noel:1967:SCA

[375] Y. Noel. Siderurgie et calcul analogique. Mathematics and Computers in Simulation, 9(4):172–173, October 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800119.

Vroman:1967:C

[376] J. Vroman. Correspondance. Mathematics and Computers in Simulation, 9(4):174–175, October 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800120.

Santesmases:1967:C

[377] J. Santesmases. Correspondance. Mathematics and Computers in Simulation, 9(4):175, October 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800132.

Maybach:1967:HAD

[378] R. L. Maybach. Hybrid analogdigital sample-averaging computer for ASTRAC II. Mathematics and Computers in Simulation, 9(4):176– 181, October 1967. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800144.

Alais:1967:MGC

[379] Pierre Alais. Méthode générale de calcul de systèmes linéaires en oscillation forcée à partir de considérations énergétiques. (French) [general method of calculating linear system in oscillation force]. Mathematics and Computers in Simulation, 9(4):182–187, October 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800156.

Bingulac:1967:HAA

[380] S. P. Bingulac, M. M. Milovanovic, B. J. Lazarevic, and B. Lolic. A hybrid approach to automatic rescaling on general purpose analog computers. Mathematics and Computers in Simulation, 9(4):188–194, October 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800168.

Degreef:1967:CTE

[381] J. F. Degreef. Cylindrical thermal equations for the analog computer. Mathematics and Computers in Simulation, 9(4):195–197, October 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546780017X.

Kokinopoulos:1967:DADb

[382] E. F. Kokinopoulos. Dynamic aseismic design of an elevated water tower. Mathematics and Computers in Simulation, 9(4):197, October 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800181.

Anonymous:1967:RAd

[383] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 9(4):198–199, October 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475467800193.

Anonymous:1967:CBd

[384] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 9(4):200–207, October 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546780020X.

Anonymous:1967:DRd

[385] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 9(4):208–215, October 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800211.

Anonymous:1967:Id

[386] Anonymous. Informations. Mathematics and Computers in Simulation, 9 (4):216–221, October 1967. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800223.

Anonymous:1967:ALc

[387] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 9(4):222–223, October 1967. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475467800235.

Anonymous:1968:PJa

[388] Anonymous. Pages 2–64 (January 1968).

Mathematics and Computers in Simulation, 10(1):??, January 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Hoffmann:1968:PNR

[389] J. Hoffmann. Présentation nouvelle de la revue. Mathematics and Computers in Simulation, 10(1):2, January 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800230.

Hoffmann:1968:NAR

[390] J. Hoffmann. New appearance of the review. Mathematics and Computers in Simulation, 10(1):3, January 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800242.

Cuenod:1968:RCF

[391] M. Cuenod. Results and conclusions of the fifth AICA meeting. Mathematics and Computers in Simulation, 10(1):4–5, January 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800254.

Carling:1968:HCS

[392] L. N. Carling. Hybrid computer solution of heat exchanger partial differential equations. Mathematics and Computers in Simulation, 10 (1):6–18, January 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800266.

Conant:1968:MDC

[393] Brian K. Conant. Modern digital control for a small iterative analog computer: the APE II digital module. Mathematics and Computers in Simulation, 10(1):19–23, January 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800278.

Santesmases:1968:C

[394] J. G. Santesmases. Correspondance.

Mathematics and Computers in Simulation, 10(1):25, January 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546880028X.

Santesmases:1968:ARM

[395] J. G. Santesmases. Authors reply Mr. Vroman’s letter. Mathematics and Computers in Simulation, 10 (1):25–26, January 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800291.

Doustens:1968:PER

[396] A. Doustens, J. Piquemal, and C. Thirriot. Procede d’exploitation rapide des reseaux a resistances. Mathematics and Computers in Simulation, 10(1):27–28, January 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800308.

Bajenesco:1968:MSD

[397] Titu Bajenesco. Un multiplicateur simple a diodes. Mathematics and Computers in Simulation, 10 (1):29–31, January 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546880031X.

Anonymous:1968:RAa

[398] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 10(1):32, January 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800321.

Anonymous:1968:CBa

[399] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 10(1):33–44, January 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800333.

Anonymous:1968:DRa

[400] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 10(1):45–54, January 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800345.

Anonymous:1968:Ia

[401] Anonymous. Informations. Mathematics and Computers in Simulation, 10(1):55–60, January 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800357.

Anonymous:1968:ALa

[402] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 10(1):61–64, January 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800369.

Anonymous:1968:PA

[403] Anonymous. Pages 64–128 (April 1968).

Mathematics and Computers in Simulation, 10(2):??, April 1968. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Brennan:1968:E

[404] Robert D. Brennan. Editorial. Mathematics and Computers in Simulation, 10(2):64, April 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800011.

Radu:1968:RAC

[405] A. Radu. Representation analogique des champs thermiques et de diffusion four l’etude de la condensation dans les elements de construction. Mathematics and Computers in Simulation, 10(2):65–71, April 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800023.

Kuntze:1968:LOE

[406] Karlheinz Kuntze. Die Lösung von Optimierungsaufgaben auf einem elektronischen Analogrechner mit Hilfe der Gradientenmethode. (German) [The solution of optimization tasks on an electronic analog calculator using the gradient method]. Mathematics and Computers in Simulation, 10(2):72–75, April 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800035.

Kingma:1968:DMD

[407] Y. J. Kingma. Design method for digital dead beat controllers. Mathematics and Computers in Simulation, 10(2):76–79, April 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800047.

Mosca:1968:TBN

[408] V. Mosca. Transient behaviour of nonminimum phase-shift networks. Mathematics and Computers in Simulation, 10(2):80–84, April 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800059.

Keen:1968:IIA

[409] Arthur W. Keen and Jacqueline L. Glover. An immittance invertor approach to RC oscillator realization using operational amplifiers. Mathematics and Computers in Simulation, 10(2):85–86, April 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800060.

Anonymous:1968:RAb

[410] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 10(2):87, April 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800072.

Anonymous:1968:CBb

[411] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 10(2):88–112, April 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800084.

Anonymous:1968:DRb

[412] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 10(2):113–121, April 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800096.

Anonymous:1968:Ib

[413] Anonymous. Informations. Mathematics and Computers in Simulation, 10(2):122–125, April 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800102.

Anonymous:1968:ALb

[414] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 10(2):126–128, April 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800114.

Anonymous:1968:PJb

[415] Anonymous. Pages 130–185 (July 1968).

Mathematics and Computers in Simulation, 10(3):??, July 1968. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ricard:1968:S

[416] J. H. Ricard. La Simulation. Mathematics and Computers in Simulation, 10(3):130–131, July 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800126.

Kuntze:1968:LOZ

[417] Karlheinz Kuntze. Die Lösung von Optimierungsaufgaben mit Zeitabhaengigen Parametern auf einem Elektronischen Analogrechner. (German) [The solution of optimization tasks with time-dependent parameters on an electronic analog computer]. Mathematics and Computers in Simulation, 10 (3):132–134, July 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800138.

Honnell:1968:GAA

[418] P. M. Honnell and R. J. Mulholland. Generation of angle and amplitude modulation functions on the matric computor. Mathematics and Computers in Simulation, 10 (3):135–143, July 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546880014X.

Ignatiev:1968:CCA

[419] M. B. Ignatiev. The checking and correction of the analog and hybrid computation by the redundant variables method. Mathematics and Computers in Simulation, 10(3):144–150, July 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800151.

Kontos:1968:HCU

[420] John Kontos. Hybrid computing units using a pulse ratio modulator. Mathematics and Computers in Simulation, 10(3):151–154, July 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800163.

Pillmeier:1968:ROM

[421] Rudolf J. Pillmeier and Joseph H. Dial.

Relay operated memory-pair. Mathematics and Computers in Simulation, 10(3):155–158, July 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800175.

Anonymous:1968:RAc

[422] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 10(3):159–160, July 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800187.

Anonymous:1968:CBc

[423] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 10(3):161–171, July 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800199.

Anonymous:1968:DRc

[424] Anonymous. Documents Reçus. Mathematics and Computers in Simulation, 10(3):172–177, July 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800205.

Anonymous:1968:Ic

[425] Anonymous. Informations. Mathematics and Computers in Simulation, 10(3):178–183, July 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800217.

Anonymous:1968:AA

[426] Anonymous. Activité de l’association. Mathematics and Computers in Simulation, 10(3):184–185, July 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800229.

Anonymous:1968:PO

[427] Anonymous. Pages 188–244 (October 1968). Mathematics and Computers in Simulation, 10(4):??, October 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Fossard:1968:E

[428] A. J. Fossard. Editorial. Mathematics and Computers in Simulation, 10 (4):188–189, October 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800370.

Smith:1968:ACG

[429] D. A. Smith and A. C. Soudack. Analog computer gradient techniques for parameter optimization. Mathematics and Computers in Simulation, 10 (4):190–198, October 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800382.

Ung:1968:UPL

[430] Man T. Ung and Gerald Klein. The use of parallel logic in teaching engineering and science. Mathematics and Computers in Simulation, 10 (4):199–210, October 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800394.

Popov:1968:QSO

[431] V. M. Popov, N. N. Tataru, and T. I.

Bajenesco. Quelques solutions originales utilisees dans la construction d’une grande calculatrice analogique. Mathematics and Computers in Simulation, 10(4):211–213, October 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800400.

Anonymous:1968:RAd

[432] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 10(4):214, October 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475468800412.

Anonymous:1968:CBd

[433] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 10(4):215–226, October 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800424.

Anonymous:1968:DRd

[434] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 10(4):227–236, October 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800436.

Anonymous:1968:Id

[435] Anonymous. Informations. Mathematics and Computers in Simulation, 10 (4):237–240, October 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475468800448.

Anonymous:1968:ALc

[436] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 10(4):240–242, October 1968. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847546880045X.

Anonymous:1968:IS

[437] Anonymous. Index — sommaire 1967– 1968. Mathematics and Computers in Simulation, 10(4):243–244, October 1968. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475468800461.

Anonymous:1969:PJa

[438] Anonymous. Pages 2–57 (January 1969).

Mathematics and Computers in Simulation, 11(1):??, January 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Weill:1969:SIC

[439] Jacky Weill. Symposium international sur le calcul analogique et hybride appliqué à l’énergie nucléaire (Versailles, 16–17–18 septembre 1968). (French) [International symposium on analog and hybrid calculation applied to nuclear energy (Versailles, 16–17–18 September 1968)]. Mathematics and Computers in Simulation, 11(1):2–3, January 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800017.

Missen:1969:SHT

[440] J. I. Missen and D. W. Stops. The solution of heat transfer problems by automated Liebmann networks. Mathematics and Computers in Simulation, 11(1):4–10, January 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800029.

Dekker:1969:STP

[441] Ir. L. Dekker. Solving two-point boundary-value problems by a levelcontrol method. Mathematics and Computers in Simulation, 11(1):11– 20, January 1969. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800030.

Fossard:1969:SFS

[442] A. J. Fossard and J. J. Eltgen. Simulation et formulation des systemes multidimensionnels. Mathematics and Computers in Simulation, 11 (1):21–23, January 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800042.

Anonymous:1969:RAa

[443] Anonymous. Résumés d’articles. Mathematics and Computers in Simulation, 11(1):27, January 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800054.

Anonymous:1969:CBa

[444] Anonymous. Chronique bibliographique. Mathematics and Computers in Simulation, 11(1):28–36, January 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800066.

Anonymous:1969:DRa

[445] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 11(1):37–45, January 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800078.

Anonymous:1969:Ia

[446] Anonymous. Informations. Mathematics and Computers in Simulation, 11(1):46–54, January 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546980008X.

Anonymous:1969:ALa

[447] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 11(1):55–57, January 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800091.

Anonymous:1969:PA

[448] Anonymous. Pages 60–120 (April 1969).

Mathematics and Computers in Simulation, 11(2):??, April 1969. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Yastreboff:1969:TOC

[449] M. Yastreboff. Time optimal control of time varying third order plants. Mathematics and Computers in Simulation, 11(2):60–64, April 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800108.

Apter:1969:ERP

[450] H. Apter. Exemple de resolution par voie analogique de probleme d’optimalisation. Mathematics and Computers in Simulation, 11(2):65–74, April 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546980011X.

Maria-Sube:1969:TPT

[451] Roland Maria-Sube. Traitement par des techniques analogiques et digitales associées d’un problème d’aèrodynamique non linéaire comportant des frontières libres. (French) [treatment with analog and digital techniques associated with a nonlinear aerodynamic problem with free borders]. Mathematics and Computers in Simulation, 11(2):75–85, April 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800121.

Belt:1969:SCD

[452] J. E. Belt and Cliff A. Harris. Simulation of a classical detection problem: Detection of a phase-coherent rf pulse train of unknown phase in additive Gaussian noise. Mathematics and Computers in Simulation, 11(2):86–89, April 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800133.

Anonymous:1969:RAb

[453] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 11(2):90–91, April 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800145.

Anonymous:1969:CBb

[454] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 11(2):92–102, April 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800157.

Anonymous:1969:DRb

[455] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 11(2):103–114, April 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800169.

Anonymous:1969:Ib

[456] Anonymous. Informations. Mathematics and Computers in Simulation, 11(2):115–118, April 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800170.

Anonymous:1969:ALb

[457] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 11(2):119–120, April 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800182.

Anonymous:1969:PJb

[458] Anonymous. Pages 122–182 (July 1969).

Mathematics and Computers in Simulation, 11(3):??, July 1969. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Richalet:1969:IS

[459] J. Richalet. Identification et simulation. Mathematics and Computers in Simulation, 11(3):122–123, July 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800194.

Trevor:1969:DLD

[460] A. Trevor and J. V. Wait. Diffe: an on-line differential-equation-solving routine with automatically scaled display. Mathematics and Computers in Simulation, 11(3):124–130, July 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800200.

Andre:1969:EQR

[461] P. Andre and J.-J. Gagnepain. Étude de quelques régimes transitores d’une diode tunnel sur le calculateur analogique à courants porteurs ANALAC A 110. (French) [study of some transitory regimes of a tunnel diode on the analog computer ANALAC A 110 with power lines]. Mathematics and Computers in Simulation, 11 (3):131–137, July 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800212.

Pakleza:1969:SME

[462] Jaroslaw Pakleza. Un simulateur modulaire d’echanges thermiques transitoires. Mathematics and Computers in Simulation, 11(3):138–141, July 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800224.

Loc:1969:SAE

[463] T. A. Phuoc Loc. Simulation analogique d’équations de diffusion non linéaires en régime transitoire. (French) [analog simulation of nonlinear diffusion equations in transitional regime]. Mathematics and Computers in Simulation, 11(3):142–149, July 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800236.

Efthymiatos:1969:OST

[464] D. Efthymiatos and W. Arndt. Sur l’optimisation des systémes en temps réel. (French) [On the optimization of systems in real time]. Mathematics and Computers in Simulation, 11(3):150–156, July 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800248.

Anonymous:1969:RAc

[465] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 11(3):157–158, July 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546980025X.

Anonymous:1969:CBc

[466] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 11(3):159–166, July 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800261.

Anonymous:1969:DRc

[467] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 11(3):167–172, July 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800273.

Anonymous:1969:Ic

[468] Anonymous. Informations. Mathematics and Computers in Simulation, 11(3):173–179, July 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800285.

Anonymous:1969:ALc

[469] Anonymous. Activité de l’Association. Mathematics and Computers in Simulation, 11(3):180–182, July 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800297.

Anonymous:1969:PO

[470] Anonymous. Pages 184–253 (October 1969). Mathematics and Computers in Simulation, 11(4):??, October 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

M:1969:LTT

[471] Wayne L. Y. M. On-line tensile test data reduction using a remote analog computer. Mathematics and Computers in Simulation, 11 (4):184–188, October 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800303.

Nougaret:1969:DAO

[472] M. Nougaret and S. Giraud. Determination analogique de l’observabilite et de la commandabilite des systems echantillonnes. Mathematics and Computers in Simulation, 11(4): 189–194, October 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800315.

Danthine:1969:SCA

[473] A. Danthine and C. Van Den Hove.

Simulation de la commande automatique de positionnement des cylindres de laminoirs. Mathematics and Computers in Simulation, 11(4): 195–204, October 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800327.

DeCarli:1969:HMI

[474] A. De Carli and T. Raimondi. Hybrid model of an inventer-induction motor system. Mathematics and Computers in Simulation, 11(4):205– 211, October 1969. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800339.

Renard:1969:CHR

[475] G. Renard and J. Quinio. Le C.H.R. 10.000 calculateur hybride à reseaux à commutation électronique. (French) [the C.H.R. 10,000 hybrid calculator for electronic switching]. Mathematics and Computers in Simulation, 11 (4):212–225, October 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800340.

Vansteenkiste:1969:SDE

[476] Ghislain C. Vansteenkiste. Study of the dynamic errors introduced by the delays inherent on combined analogdigital computation in hybrid systems. Mathematics and Computers in Simulation, 11(4):226–232, October 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800352.

Anonymous:1969:RAd

[477] Anonymous. Résumés d’articlse. Mathematics and Computers in Simulation, 11(4):233–235, October 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800364.

Anonymous:1969:E

[478] Anonymous.  Errata.  Mathemat-

ics and Computers in Simulation, 11 (4):235, October 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475469800376.

Anonymous:1969:CBd

[479] Anonymous. Chronique bibliographique.

Mathematics and Computers in Simulation, 11(4):236–243, October 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800388.

Anonymous:1969:DRd

[480] Anonymous. Documents reçus. Mathematics and Computers in Simulation, 11(4):244–248, October 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847546980039X.

Anonymous:1969:Id

[481] Anonymous. Informations. Mathematics and Computers in Simulation, 11 (4):249–251, October 1969. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800406.

Anonymous:1969:ALd

[482] Anonymous. Activité de l’Association.

Mathematics and Computers in Simulation, 11(4):252–253, October 1969. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475469800418.